Identify the vertex of the graph of f(x)=(x-5)^2-8
Answer by Math_prodigy(24)

Answer: vertex is at (5,-8)
f(x)= X^2 is parabola with the vertex at (0,0)
f(x)=(x-5)^2-8 is transformed parabola. 5 units to the right and 8 units down.
(0+5,0-8) =(5,-8)